A ball is dropped off a high cliff, and 2 s later another ball is thrown vertically downward 

with an initial speed of 30 m/s. How long will it take the second ball to overtake the first?


1
Expert's answer
2021-09-17T09:23:09-0400

Let's first write the vertical displacement of the first and second balls:


"y_1=\dfrac{1}{2}gt^2,""y_2=v_0(t-2)+\dfrac{1}{2}g(t-2)^2."

When the second ball overtakes the first one "y_1=y_2" and we can write:


"\dfrac{1}{2}gt^2=v_0(t-2)+\dfrac{1}{2}g(t-2)^2,""\dfrac{1}{2}gt^2=v_0t-2v_0+\dfrac{1}{2}gt^2-2gt+2g,""30t-60-19.6t+19.6=0,""10.4t-40.4=0,""t=\dfrac{40.4}{10.4}=3.88\ s."